# Securing Large Language Models: Threats, Vulnerabilities and Responsible Practices

### v1

@@ 6.2 Red/Green Teaming
For instance, Zhuo et al. Zhuo et al. (2023), have investigated whether ChatGPT produces hazardous outputs by employing prompt-injection techniques, while other studies Shi et al. (2023b); Casper et al. (2023); Perez et al. (2022a), focus on specific red-teaming aspects including developing toxicity classifiers or using red-teaming LLMs to identify risky generations that might otherwise go unnoticed.
A work by Ganguli et al. Ganguli et al. (2022) demonstrates how various sampling techniques could discourage particular red-teaming elements. RLHF, for instance, is shown to be more resilient compared to rejection sampling. However, these approaches often involve substantial human involvement.
To address this manual burden, other researchers have sought ways to automate red teaming. Lee et al. Lee et al. (2023a), for example, employ Bayesian optimization to conduct red teaming with minimal queries and reduced reliance on human assistance.
@@ 6.3.1 Fine-tuning Language Models as Supervised Detectors
Another drawback lies in their sensitivity to paraphrasing attacks. These attacks involve adding a paraphrase on top of the generative text model, which can deceive any form of detector, including those utilizing supervised neural networks.
@@ 6.3.2 Pre-trained Language Models as Zero-shot Detectors
According to Mitchell et al. (2023), these techniques often set a threshold for the predicted per-token log probability to identify AI-generated texts. This approach relies on the observation that passages generated by AI often exhibit a negative log probability curvature.
@@ 6.3.3 Detection based on Information Retrieval Techniques
These techniques operate by comparing a given text with a database of texts generated by LLMs. The objective is to identify semantically similar matches, thereby aiding in the discrimination process.
By leveraging these approaches, researchers aim to enhance the robustness and reliability of text detection mechanisms. Whether it’s matching keywords, traversing hypertext links, or employing more sophisticated algorithms, the goal remains consistent: to discern the subtle nuances that distinguish human-authored content from its AI-generated counterparts Krishna et al. (2023); Sadasivan et al. (2023).
@@ 6.3.4 Watermarking as a Signature for Detection
Another work by Kirchenbauer et al. Kirchenbauer et al. (2023b) examines the reliability of watermarks as a method for identifying and keeping track of AI-generated text. They investigate how well watermarked text holds up against human restructuring, non-watermarked LLM paraphrasing, and blending into lengthier human-written documents.
@@
Papers Method Main Idea Vulnerabilities Solaiman et al. (2019); Bakhtin et al. (2019); Antoun et al. (2023); Zhan et al. (2023); Li et al. (2023b) Supervised detection To fine-tune a model on sets of AI and human generated texts. Training on commonly used datasets, makes it vulnerable to most attacks including paraphrasing.  Su et al. (2023); Zer (2023); Wang et al. (2023b); Gehrmann et al. (2019) Zero-shot detection To use a pre-trained language model in zero-shot settings. Reduces the risk of data poisoning attacks and eliminates data and resource over- heads, but it is still susceptible to other adversarial attacks like spoofing and paraphrasing. Krishna et al. (2023); Sadasivan et al. (2023) Retrival-based detection Apply methods of information retrieval to match a given text with a collection of texts generated by LLMs and finding similarities in meaning. It is impractical because it requires a large and updated collection of texts, which is computationally expensive, or may be unavailable for all domains, tasks or models. It is also vulnerable to paraphrasing and spoofing attacks. Kirchenbauer et al. (2023a); Yang et al. (2023b); Kirchenbauer et al. (2023b); Sadasivan et al. (2023) Watermarking To use a model signature in the produced text outputs to stamp particular pattern. The most trustworthy strategy, but is shown to be fundamentally impossible for generative models. It is susceptible to attacks such as rewording and spoofing. Yu et al. (2023); Yang et al. (2023c); Mitchell et al. (2023); Su et al. (2023) Feature-based detection To identify and classify based on extracted discriminating features. Susceptible to adversarial attacks such as paraphrasing.
@@ 6.3.5 Discriminating Features as Detection Clues
Another stream of work is to identify and classify based on discriminating features.
For instance, Yu et al. Yu et al. (2023) have identified a genetic inheritance characteristic specific to GPT-generated text. According to this characteristic, the model’s output essentially rearranges the content present in its training corpus. In simpler terms, when repeatedly answering a question, the model’s responses remain constrained by the information within its training data, resulting in limited variations. This hypothesis suggests that the output of a language model (such as ChatGPT) is predictable, implying that for highly similar questions, the model will produce correspondingly similar answers. Analogously, paternity testing involves using DNA profiles to determine whether an individual is the biological parent of another individual. This process becomes particularly crucial when parental rights and responsibilities are in question, and uncertainty exists regarding a child’s paternity.
@@
Another discriminating feature is the susceptibility of text to manipulations. AI-generated and human-written texts are both negatively affected by small perturbations e.g., replacing some of the words. However, some recent work Mitchell et al. (2023); Su et al. (2023) reveal that AI-generated text is more susceptible to such manipulations.
For instance, to measure sensitivity of LLMs to perturbations, Su et al. propose Log-Likelihood Log-Rank Ratio (LRR):
@@ 6.3.6 Generalizability of Detection Techniques
There are few works such as a study by Wang et al. Wang et al. (2023h) that investigate the generalizability of detectors by conducting experiments on a large-scale corpus spanning multiple generators, domains, and languages. Their investigation involves leveraging various generative models, including ChatGPT, textdavinci-003, LLaMa, FlanT5, Co-here, Dolly-v2, and BLOOMz, to create text articles. Subsequently, they attempt to distinguish between AI-generated and human-written content using both traditional machine learning methods (e.g., Linear Support Vector Machine) and modern transformer-based models (e.g., RoBERTa), focusing on stylistic features. Interestingly, their findings reveal that while these text detection methods perform well within their specific domains, they encounter challenges in out-of-domain detection tasks.
In addition, they discover that all detection models perform better for detecting content that exhibits a particular pattern which sets it apart from content written by human (ChatGPT in this case).
@@ 6.3.7 Vulnerabilities of Detection Techniques
As previously mentioned, zero-shot attacks are susceptible to adversarial techniques like data poisoning. Researchers employ supervised methods to counter these attacks, but most detection strategies still remain vulnerable to paraphrasing or spoofing.
@@
Therefore, watermarking generative models may be fundamentally unachievable, necessitating alternative approaches to protect the intellectual property of model developers and LLM users. Table 6 demonstrates an overview of detection strategies, highlighting the vulnerabilities associated with each category.
@@ 6.3.8 Discussion on the Possibility of Detection
Sadasivan et al. for instance, come up with an impossibility finding Sadasivan et al. (2023): “as language models become more sophisticated and better at emulating human text, the performance of even the best-possible detector decreases drastically”. Sadasivan et al. propose an upper bound for the area under the ROC curve of any decoder $\mathcal{D}$ as:
@@
where $TV(\mathcal{M},\mathcal{H})$ is the total variation distance between machine and human generated texts. This formula indicates that when human and machine generated texts are very similar i.e., $TV(\mathcal{M},\mathcal{H})$ is very small, even the best-possible detector may only perform marginally better than a random classifier.
@@
7 New Opportunities and Future Research
@@
7.1 Opportunities in Security & privacy Study
Opportunities in Data Leakage and Memorization
@@ 7.1 Opportunities in Security & privacy Study
  • Developing multi-aspect techniques to prevent sensitive data leakage: These techniques should consider various dimensions, including aspects related to the model (such as training data selection and differential privacy), data (such as data classification, access control, and monitoring), and user-based factors (such as detecting abnormal patterns in user-LLM interactions and managing user access)
  • Developing new methods to mitigate memorization: Given the current scarcity of techniques for tackling memorization in LLMs, it is crucial to put forth novel approaches—such as model editing—to mitigate the memorization.
@@ 7.2 Opportunities in Vulnerabilities Study
  • Extending experiments on diverse NLP applications besides classification task: While classification tasks have been the primary focus, exploring other NLP applications—such as language generation, summarization, sentiment analysis, and question answering—will provide a more comprehensive understanding of LLM behavior and security implications.
@@
  • Adapting multi-faceted defense strategy for safeguarding against backdoor injection: Researchers have the opportunity to craft multi-faceted defense strategies against backdoor injection by combining techniques such as prompt filtering mechanisms to exclude harmful input and specialized security-oriented LLMs trained to detect suspicious instructions.
@@ 7.3 Opportunities in Risk Mitigation Study
AI Generated Text Detection Techniques
As mentioned earlier, the detection of AI-generated text is a crucial and challenging task and the current methods are frequently constrained by different factors, and susceptible to malicious attacks. Therefore, it is critical to pursue further exploration and investigation on both theoretical and practical aspects of AI-text detection. Some of the opportunities are:
  • Creating more diverse and representative datasets: The existing datasets may not cover all the nuances of AI-generated content used for training and evaluating AI-text detection models. Developing more diverse and representative datasets improves the models’ ability to generalize and enables more reliable assessments.
  • Exploring more advanced and interpretable features: By discerning subtle nuances and interpretable features, we can gain a fine-grained understanding of the distinctions between human-written and AI-generated texts.
@@
  • A comprehensive understanding of the fundamental feasibility and boundaries: A thorough grasp of the fundamental feasibility and limitations of AI-generated text detection is essential. However, the theoretical aspect of this task is mainly overlooked in current state-of-the-art methods. Therefore, further exploration and investigation of the theoretical aspects are necessary. This can lead to the development of more robust and effective techniques as well as uncovering new research directions.
@@
Similar to the cyber-security domain, where red teaming has been beneficial for enhancing security, both red teaming and green teaming for LLMs have revealed the vulnerability of LLMs to malicious users. There have been some nascent contributions like RLHF that can prevent more attacks than other methods, as demonstrated in Perez et al. (2022a). However, there are still some areas for improvement in these areas such as:
  • Creating more safeguards to prevent the attacks affecting the LLMs: Given the rising popularity of LLMs, they are becoming the central component in many products. Thus, there is a critical need to implement multiple safeguards, as new attacks continually emerge.
@@ 8 Conclusions
This paper provides a comprehensive analysis of the security and risk mitigation aspects of LLMs. We examine the security issues that emerge with LLM usage, such as information leakage, unauthorized access, and insecure code generation. In addition, we explore various types of attacks that target LLMs and classify them into three main categories: Model-based, training-time and inference-time attacks. We also investigate the risks and misuses of LLMs, such as bias, discrimination, misinformation, plagiarism, copyright infringement and other potential social and ethical implications of applying LLMs in different domains. Moreover, we present a thorough evaluation of the mitigation strategies that can be employed to improve the security and robustness of LLMs, such as red and green teaming, model editing, watermarking, and AI-generated text detection techniques, while discussing the limitations and trade-offs of each strategy. Lastly, we identify some open challenges and future directions for research in this area, such as developing more effective and efficient defense mechanisms, establishing standards and guidelines for LLM development and deployment, promoting collaboration and awareness among the stakeholders involved in LLM utilization.
@@ 9 Acknowledgements
The authors express their gratitude to Dr. Sadid Hassan for engaging in several discussions and providing valuable feedback on the paper’s content. This study represents independent research conducted by the authors and does not necessarily represent the views or opinions of any organizations.

### v2

@@ 6.2 Color Teaming
For instance, Zhuo et al. Zhuo et al. (2023), investigate whether ChatGPT produces hazardous outputs by employing prompt-injection methods, while other studies Shi et al. (2023b); Casper et al. (2023); Perez et al. (2022a), focus on specific red-teaming aspects including developing toxicity classifiers or identifying risky generations that might otherwise go unnoticed.
A work by Ganguli et al. Ganguli et al. (2022) explores how various sampling techniques could discourage particular red-teaming elements. RLHF, for instance, is shown to be more resilient compared to rejection sampling. However, these approaches often involve substantial human involvement. To address this manual burden, other researchers have sought ways to automate red teaming. Lee et al. Lee et al. (2023a), for example, employ Bayesian optimization to conduct red teaming with minimal queries and reduced reliance on human assistance.
@@
Other Teaming Terms
Beyond the red–green distinction, researchers have recently begun exploring purple, blue, and even “rainbow” teaming, where the goal is to combine or extend adversarial discovery with in-situ hardening of the model.
Purple teaming marries the attacker’s perspective with immediate, automated defenses. Zhou et al. Zhou et al. (2024a) introduce Purple-teaming LLMs with Adversarial Defender training (PAD), a self-play pipeline in which an attacker LLM continuously elicits unsafe responses while a defender LLM learns—GAN-style—to detect and rebut them, markedly improving safety without sacrificing utility. Complementing PAD, Purple Llama CyberSecEval offers a large-scale benchmark that probes coding assistants for two concrete risks—suggesting insecure code and complying with cyber-attack requests—thereby providing a quantitative basis for purple-team evaluation Bhatt et al. (2023).
On the blue-team, Zhao et al. Zhao et al. (2025) propose BlueSuffix, a reinforced suffix generator coupled with visual + textual purifiers that harden vision–language models against multi-modal jailbreaks while preserving benign performance.
Finally, Rainbow Teaming frames adversarial-prompt discovery as an open-ended quality–diversity search, yielding hundreds of transferable attacks and demonstrating that fine-tuning on this diverse corpus can simultaneously raise safety and maintain helpfulness Samvelyan et al. (2024).
@@ 6.3.1 Fine-tuning Language Models as Supervised Detectors
Another drawback lies in their sensitivity to paraphrasing attacks. These attacks often add a paraphraser on top of a generative model, which can deceive any form of detector, including those utilizing supervised neural networks.
@@ 6.3.2 Pre-trained Language Models as Zero-shot Detectors
According to Mitchell et al. (2023), these techniques often set a threshold for the predicted per-token log probability to identify AI-generated texts. This approach relies on the observation that passages generated by AI often exhibit a negative log probability curvature. Specifically, AI-generated text, $x\sim p_{\theta}(\cdot)$, tends to lie in regions of negative curvature in the log-likelihood landscape, $\log p(x)$, where nearby samples (i.e., similar texts) generally have lower model log-probability on average. In contrast, human-written text, $x\sim p_{\text{real}}(\cdot)$, typically does not occupy regions with clearly defined negative curvature, and nearby samples may exhibit either higher or lower log-probability without a consistent pattern as shown in Figure 9.
Figure 9: Log-likelihood curvature for AI vs. human-generated text. This figure is taken from Mitchell et al. (2023)
@@ 6.3.3 Detection based on Information Retrieval Techniques
These techniques operate by comparing a given text with a database of texts generated by LLMs. The objective is to identify semantically similar matches, thereby aiding in the discrimination process. By leveraging these approaches, researchers aim to enhance the robustness and reliability of text detection mechanisms. Whether it’s matching keywords, traversing hypertext links, or employing more sophisticated algorithms, the goal remains consistent: to discern the subtle nuances that distinguish human-authored content from its AI-generated counterparts Krishna et al. (2023); Sadasivan et al. (2023).
@@ 6.3.4 Watermarking as a Signature for Detection
Kirchenbauer et al. Kirchenbauer et al. (2023b) examine the reliability of watermarks as a method for identifying and keeping track of AI-generated text. They investigate how well watermarked text holds up against human restructuring, non-watermarked LLM paraphrasing, and blending into lengthier human-written documents.
@@
{tblr} colspec = p2.4cmp1.4cmp3.5cmp4.8cm, row1 = bg=orange!20, row2 = bg=orange!2, row3 = bg=orange!2, row4 = bg=orange!2, row5 = bg=orange!2, row6 = bg=orange!2, PapersMethodMain IdeaVulnerabilities
Solaiman et al. (2019); Bakhtin et al. (2019); Antoun et al. (2023); Zhan et al. (2023); Li et al. (2023b)Supervised detectionTo fine-tune a model on sets of AI and human generated texts. Training on commonly used datasets, makes it vulnerable to most attacks including paraphrasing.
 Su et al. (2023); Zer (2023); Wang et al. (2023b); Gehrmann et al. (2019)Zero-shot detectionTo use a pre-trained language model in zero-shot settings. Reduces the risk of data poisoning attacks and eliminates data and resource over- heads, but it is still susceptible to other adversarial attacks like spoofing and paraphrasing.
Krishna et al. (2023); Sadasivan et al. (2023)Retrival-based detectionApply methods of information retrieval to match a given text with a collection of texts generated by LLMs and finding similarities in meaning. It is impractical because it requires a large and updated collection of texts, which is computationally expensive, or may be unavailable for all domains, tasks or models. It is also vulnerable to paraphrasing and spoofing attacks.
Kirchenbauer et al. (2023a); Yang et al. (2023b); Kirchenbauer et al. (2023b); Sadasivan et al. (2023)WatermarkingTo use a model signature in the produced text outputs to stamp particular pattern. The most trustworthy strategy, but is shown to be fundamentally impossible for generative models. It is susceptible to attacks such as rewording and spoofing.
Yu et al. (2023); Yang et al. (2023c); Mitchell et al. (2023); Su et al. (2023) Feature-based detectionTo identify and classify based on extracted discriminating features.Susceptible to adversarial attacks such as paraphrasing.
@@ 6.3.5 Discriminating Features as Detection Clues
Another stream of work is to identify and classify based on discriminating features. For instance, Yu et al. Yu et al. (2023) have identified a genetic inheritance characteristic specific to GPT-generated text. According to this characteristic, the model’s output essentially rearranges the content present in its training corpus. In simpler terms, when repeatedly answering a question, the model’s responses remain constrained by the information within its training data, resulting in limited variations. This hypothesis suggests that the output of a language model (such as ChatGPT) is predictable, implying that for highly similar questions, the model will produce correspondingly similar answers. Analogously, paternity testing involves using DNA profiles to determine whether an individual is the biological parent of another individual. This process becomes particularly crucial when parental rights and responsibilities are in question, and uncertainty exists regarding a child’s paternity.
@@
Another discriminating feature is the susceptibility of text to manipulations. AI-generated and human-written texts are both negatively affected by small perturbations e.g., replacing some of the words. However, some recent work Mitchell et al. (2023); Su et al. (2023) reveal that AI-generated text is more susceptible to such manipulations. For instance, to measure sensitivity of LLMs to perturbations, Su et al. propose Log-Likelihood Log-Rank Ratio (LRR):
@@ 6.3.6 Generalizability of Detection Techniques
There are few works such as a study by Wang et al. Wang et al. (2023h) that investigate the generalizability of detectors by conducting experiments on a large-scale corpus spanning multiple generators, domains, and languages. Their investigation involves leveraging various generative models, including ChatGPT, textdavinci-003, LLaMa, FlanT5, Co-here, Dolly-v2, and BLOOMz, to create text articles. Subsequently, they attempt to distinguish between AI-generated and human-written content using both traditional machine learning methods (e.g., Linear Support Vector Machine) and modern transformer-based models (e.g., RoBERTa), focusing on stylistic features. Interestingly, their findings reveal that while these text detection methods perform well within their specific domains, they encounter challenges in out-of-domain detection tasks. In addition, they discover that all detection models perform better for detecting content that exhibits a particular pattern which sets it apart from content written by human (ChatGPT in this case).
@@ 6.3.7 Vulnerabilities of Detection Techniques
As mentioned earlier, zero-shot attacks are susceptible to adversarial techniques like data poisoning. Researchers employ supervised methods to counter them, but most detection strategies still remain vulnerable to paraphrasing or spoofing.
@@
Therefore, watermarking generative models may be fundamentally unachievable, necessitating alternative approaches to protect the intellectual property of model developers and LLM users. Table 6.3.4 demonstrates an overview of detection strategies, highlighting the vulnerabilities associated with each category.
@@ 6.3.8 Discussion on the Possibility of Detection
Sadasivan et al. for instance, come up with an impossibility finding Sadasivan et al. (2023): “as language models become more sophisticated and better at emulating human text, the performance of even the best-possible detector decreases drastically”. They propose an upper bound for the area under the ROC curve of any decoder $\mathcal{D}$ as:
@@
where $TV(\mathcal{M},\mathcal{H})$ is the total variation distance between machine and human generated texts. This formula indicates that when human and machine generated texts are very similar i.e., $TV(\mathcal{M},\mathcal{H})$ is very small, even the best-possible detector may only perform marginally better than a random classifier. The proof is provided in Appendix. A.
According to this formulation, as the $TV$ distance between AI and human text distributions reduces, the AUROC of the optimal detector also decreases accordingly as illustrated in Figure 10.
Figure 10: Comparing the AUROC of the optimal detector to that of a random classifier shows that as the TV distance between AI and human generated text distributions decreases, the AUROC of the optimal detector also drops accordingly.
@@
7 Opportunities and Future Research
@@
7.1 Opportunities in Security & privacy
Opportunities in Data Leakage & Memorization
@@ 7.1 Opportunities in Security & privacy
  • Developing multi-aspect techniques to prevent sensitive data leakage: These techniques should consider various dimensions, including model-related aspects (e.g., training data selection and differential privacy), data (such as data classification, access control, and monitoring), and user-based factors (e.g., detecting abnormal patterns in user-LLM interactions and managing user access)
  • Developing new methods to mitigate memorization: Given the current scarcity of techniques for tackling memorization in LLMs, it is crucial to put forth novel approaches (e.g., model editing) to mitigate the memorization.
@@ 7.2 Opportunities in Vulnerabilities Study
  • Extending experiments on diverse NLP applications besides classification task: While classification have been the primary focus, exploring other tasks (e.g., language modeling, summarization, sentiment analysis, and question answering) will provide a more comprehensive understanding of LLM behavior and security implications.
@@
  • Adapting multi-faceted defense strategy for safeguarding against backdoor injection: There is an opportunity to craft multi-faceted defense strategies against backdoor injection by combining techniques such as prompt filtering to exclude harmful input and specialized security-oriented LLMs trained to detect suspicious instructions.
@@ 7.3 Opportunities in Risk Mitigation Study
AI-Generated Text Detection Techniques
As mentioned earlier, the detection of AI-generated text is a challenging task and the current methods are frequently constrained by different factors, and susceptible to malicious attacks. Therefore, it is critical to pursue further exploration and investigation on both theoretical and practical aspects of AI-text detection. Some of the opportunities are:
  • Creating diverse and representative datasets: The existing datasets may not cover all the nuances of AI-generated content used to train and evaluate AI-text detection models. Developing more diverse and representative datasets improves the models’ ability to generalize and enables more reliable assessments.
  • Exploring more advanced and interpretable features: By discerning subtle nuances and interpretable features, we can gain a fine-grained understanding of the distinctions between human vs. AI-generated texts.
@@
  • A comprehensive understanding of the fundamental feasibility and boundaries: A thorough grasp of the fundamental feasibility and limitations of AI-generated text detection is essential. However, the theoretical aspect of this task is mainly overlooked in the literature. Therefore, further exploration and investigation of the theoretical aspects are necessary. This can lead to the development of more robust and effective techniques as well as uncovering new research directions.
@@
Similar to the cyber-security domain, where red teaming has been beneficial to enhance security, both red teaming and green teaming for LLMs have revealed the vulnerability of LLMs to malicious users. There are nascent contributions like RLHF that prevent more attacks than other methods, as demonstrated in Perez et al. (2022a). However, there are still some areas for improvement such as:
  • Creating more safeguards to prevent the attacks: Given the rising popularity of LLMs, they are becoming the central component in many products. Thus, there is a critical need to implement multiple safeguards, as new attacks continually emerge.
@@ 8 Conclusions
This paper provides a comprehensive analysis of the security and risk mitigation aspects of LLMs. We examine the security issues that emerge with LLM usage, such as information leakage, unauthorized access, and insecure code generation. In addition, we explore various types of attacks that target LLMs and classify them into three main categories: Model-based, training-time and inference-time attacks. We also investigate the risks and misuses of LLMs, such as bias, discrimination, misinformation, plagiarism, copyright infringement and other potential social and ethical implications of applying LLMs in different domains. Moreover, we present a thorough evaluation of the mitigation strategies that can be employed to improve the security and robustness of LLMs, such as red and green teaming, model editing, watermarking, and AI-generated text detection techniques, while discussing the limitations and trade-offs of each strategy. Lastly, we identify some open challenges and future directions for research in this area, such as developing more effective defense mechanisms, establishing standards and guidelines for development and deployment of LLMs, promoting collaboration and awareness among the stakeholders involved in utilization.
@@ 9 Acknowledgements
The authors express their gratitude to Dr. Sadid Hassan for engaging in several discussions and providing valuable feedback on this work. This study represents independent research conducted by the authors and does not necessarily represent the views or opinions of any organizations.
Appendix A Appendix
A-ROC and AUROC Bound using Total Variation
The ROC curve is a plot between the true positive rate (TPR) and the false positive rate (FPR), which are defined as follows Sadasivan et al. (2023):
$\text{TPR}_{\gamma}$ $=\mathbb{P}_{s\sim M}[D(s)\geq\gamma]$
$\text{FPR}_{\gamma}$ $=\mathbb{P}_{s\sim H}[D(s)\geq\gamma]$
where $\gamma$ is a classifier threshold parameter, $M$ denotes the distribution over positives (e.g., machine-generated texts), and $H$ denotes the distribution over negatives (e.g., human-written texts).
We can bound the difference between $\text{TPR}_{\gamma}$ and $\text{FPR}_{\gamma}$ by the total variation distance between distributions $M$ and $H$ Sadasivan et al. (2023):
$|\text{TPR}_{\gamma}-\text{FPR}_{\gamma}|$ $=\left|\mathbb{P}_{s\sim M}[D(s)\geq\gamma]-\mathbb{P}_{s\sim H}[D(s)\geq\gamma]\right|$
$\leq\text{TV}(M,H)$
Thus,
$\text{TPR}_{\gamma}\leq\text{FPR}_{\gamma}+\text{TV}(M,H)$
Since $\text{TPR}_{\gamma}\leq 1$, we have:
$\text{TPR}_{\gamma}\leq\min(\text{FPR}_{\gamma}+\text{TV}(M,H),1)$
Let $x=\text{FPR}_{\gamma}$, $y=\text{TPR}_{\gamma}$, and $tv=\text{TV}(M,H)$. The AUROC is defined as:
[…]
We split the integral at $x=1-tv$:
[…]
$=\left[\frac{x^{2}}{2}+tv\cdot x\right]_{0}^{1-tv}+\left[x\right]_{1-tv}^{1}$
$=\left(\frac{(1-tv)^{2}}{2}+tv(1-tv)\right)+(1-(1-tv))$
$=\frac{(1-tv)^{2}}{2}+tv(1-tv)+tv$
Simplifying the expression:
$\text{AUROC}(D)$ $\leq\frac{(1-tv)^{2}}{2}+tv(1-tv)+tv$
$=\frac{1-2tv+tv^{2}}{2}+tv-tv^{2}+tv$
$=\frac{1}{2}+tv-\frac{tv^{2}}{2}$
$\boxed{\text{AUROC}(D)\leq\frac{1}{2}+tv-\frac{tv^{2}}{2}}$